# A General-Purpose Multimodal Foundation Model for Dermatology

### v1

@@ Methods
MYM & HOP cohort (TBP). The MYM cohort [53] is an in-house dataset designed to study the natural history of melanocytic naevi. This cohort recruited individuals from the general population who were 18 years or older and had at least one mole. The HOP study [52] is an in-house sequential dataset targeting high-risk melanoma individuals. Inclusion criteria for the HOP study were: at least one melanoma (including in situ) diagnosed before the age of 40 years, two or more melanomas (including in situ) diagnosed before the age of 65, a strong family history (two or more first-degree relatives affected) and/or known pathogenic mutations in a hereditary melanoma gene, and/or a diagnosis of dysplastic naevus syndrome. For both cohorts, 3D TBP was conducted using a VECTRA WB360 (Canfield Scientific Inc., Parsippany, NJ, USA), which instantaneously captures 92 cross-polarised 2D images with standardized lighting and subsequently merges them to create a 3D avatar. The average number of lesion tiles per subject from TBP was approximately 500. Demographic factors were collected using standard questionnaires, with clinical characteristics collected by research assistants. The final dataset of automatically detected lesion image tiles $\geq$ 2 mm in diameter comprises 405,856 images. More detailed information about these two datasets can be found in our previous studies [53, 52].
@@
ACEMID pathology pilot study. The ACEMID dermatopathology dataset is an in-house collection comprising 98 slides, which have been processed to yield 80,312 patch images.
NSSI. The NSSI dataset is an in-house sequential collection containing 29,832 dermoscopic images. These images were collected as part of the Brisbane Naevus Morphology Study, conducted from approximately 2009 to 2014. The images were captured using a digital dermatoscope attached to a Fotofinder ATBM imaging system, resulting in images of approximately 0.05MB, with dimensions of 768 $\times$ 576 pixels at 96 dpi. The dataset includes up to 7 time points per participant, scheduled every 6 months over a 3-year period, with some variation due to late enrollment, scheduling non-adherence, or loss to follow-up. Individual lesions were assigned a consistent number across visits, facilitating temporal analysis.
Edu1 & Edu2. The Edu1 and Edu2 datasets are curated from in-house educational resources, such as educational notes and materials. These datasets contain 81,947 and 67,430 clinical images, respectively, providing a diverse range of examples that are typically used for training medical professionals in dermatology.
@@
Sequential data preprocessing for lesion change detection. Our proposed sequential data preprocessing method consists of dark corner removal, skin inpainting, hair removal, image registration, and lesion segmentation. For the first two steps, we follow the approach outlined in [73]. Given an image with or without dark corner artifacts (DCA), we first convert it to grayscale. We then extract the contour by applying OpenCV’s [74] binary threshold function to the grayscale image, empirically setting the threshold at 100, and using the findContours function with RETR_TREE mode and CHAIN_APPROX_SIMPLE method. We identify the largest area contour in the image, which most closely matches the edge of the DCA, by calculating the area of all existing contours. Using OpenCV’s minEnclosingCircle function, we capture a circular area that encompasses this largest contour. To mitigate the effect of gradient colors at the dark corner edges, we scale this circle down to 80% of its original radius and convert it into a binary mask. Finally, we inpaint the original image using this mask, employing OpenCV’s implementation of the Telea algorithm with an inpaint radius of 10. Following dark corner removal and inpainting, we implement a hair removal step to further improve image quality and facilitate more accurate registration. This process begins by converting the image to grayscale. We then apply a black hat morphological operation using a 17$\times$17 structuring element to isolate dark, thin structures (hairs) from the background. The resulting image is thresholded to create a binary mask of the detected hair structures. Finally, we use OpenCV’s inpaint function with the Telea algorithm to fill in the hair regions, effectively removing them from the image. This hair removal step is crucial for improving the accuracy of subsequent image registration and analysis. For image registration, we implement an AKAZE [75] feature-based approach. The process begins by detecting key points and computing descriptors using the AKAZE algorithm, which is particularly effective for non-linear scale-spaces. We utilize OpenCV’s AKAZE_create function, setting the descriptor size to 0, threshold to $9\times 10^{-5}$, and number of octaves to 4. Keypoint matching is performed using a Brute Force matcher with Hamming distance and cross-checking enabled. To refine the matches and estimate the transformation between images, we employ the RANSAC (Random Sample Consensus) algorithm implemented via skimage.measure.ransac. This estimates an EuclideanTransform model, which accounts for rotation and translation between the images. The resulting transformation is then applied to warp one image onto the other using skimage.transform.warp with reflection padding and linear interpolation.
@@
Early melanoma detection (Reader study 1). We fine-tuned our foundation model on the private SDDI-Alfred dataset[47] using a ten-fold cross-validation approach. We utilized cross-entropy loss with a learning rate of $5\times 10^{-4}$. We train models for 50 epochs with a warmup of 10 epochs. The model exhibiting the best AUROC on the validation set is selected as the final model. We then employed an out-of-fold (OOF) prediction approach to generate melanoma predictions for all sequential images. For each image sequence, we recorded the time point at which the model first made a correct diagnosis of melanoma; otherwise, the model was considered to have failed in detecting the melanoma. For the human evaluation, 12 clinicians—seven dermatologists with over five years of experience and five registrars with less than five years of experience—were invited to assess the serial dermoscopic data. The images were presented to the reviewers using Qualtrics™ (Provo, UT, USA), with the reviewers blinded to the true diagnoses. For each case, information such as the patient’s age, sex, lesion location, and date of imaging was provided. Initially, only the first dermoscopic image in the sequence was shown, and reviewers were asked to classify the lesion as either benign or malignant. As they progressed through the sequence, side-by-side image comparisons were made available to assess changes over time. Once a diagnosis was submitted, it could not be revised. To mitigate bias, we included 10 single time-point melanoma images, preventing reviewers from assuming that the first image in a series was benign. We then compared the diagnostic performance of the clinicians with our model, focusing on the time point at which a malignant diagnosis was first made by either the clinicians or the algorithm.
@@
Interaction platform, raters, reader study and implementation for human-AI collaboration. The reader study was conducted using DermaChallenge, a web-based platform developed and hosted by the Medical University of Vienna for online education on dermatoscopy, as described in a previous study [65]. To ensure proper authentication and data management, readers were required to register with a unique username, valid email address, and password. Active users on the platform, who previously actively agreed to be contacted, were recruited via a single email. Before commencing to the study phase, all users had to finish three introduction levels to be familiarized with the platforms’ user interface and image types. The number of correct answers in the first iteration of these levels normalized against the mean score of the entire DermaChallenge platform user base, served as a score of experience. Users were grouped into “Low” (n=11), “Medium” (n=21), and “High” (n=9) experience based on quantiles with cuts at 0.25 and 0.75 probability (R stats::quantile() function). Within the study level, users were shown batches of 10 images, randomly selected from a pool of 1,511 images, i.e. the ISIC 2018 Task 3 test set, with a predefined diagnosis distribution (AKIEC: 1, BCC: 1, BKL: 1, DF: 1, VASC: 1, MEL: 2, NV: 3). For each image a user had to choose one diagnosis out of seven options, and subsequently again after assistance from our foundation model, presented as multi-class probabilities visualized as bars and numbers for each class. Readers had the flexibility to complete multiple survey rounds with different image batches at their discretion, incompletely answered batches were omitted. The study was conducted online from August 20 to September 12, 2024, during which we collected data from 41 raters. Our foundation model for decision support utilized a weighted random sampler strategy, following the approach from [65] but excluding test-time augmentation. The model demonstrated robust performance, achieving an 80.4% mean (macro-averaged) recall, with notably high recall rates for critical skin lesions: 87.2% for melanoma and 86.0% for Basal Cell Carcinoma (BCC).
@@
Evaluation metrics. For multi-class tasks, we primarily use a weighted F1 score, which averages class-specific F1 scores (harmonic means of precision and recall) weighted by class size. It addresses class imbalance in multi-class scenarios. For binary classification, we primarily use AUROC (Area Under the Receiver Operating Characteristic curve), measuring the model’s ability to distinguish between classes across all classification thresholds. An AUROC of 1.0 indicates perfect classification, while 0.5 suggests random guessing. This metric is particularly useful for imbalanced datasets and when we need to evaluate trade-offs between true positive and false positive rates. For the two reader studies, we report accuracy. In skin lesion segmentation, we use the Dice Similarity Coefficient (DSC) and Jaccard index (JAC) to assess segmentation quality. For TBP-based melanoma screening, we primarily report the sensitivity (recall) in malignant lesions, focusing on the model’s ability to correctly identify malignant cases.
@@
DermNet [46] (23 classes) : DermNet contains 19,559 clinical images, the dataset consists of images of 23 types of skin diseases.
@@
MMT-74 (74 classes) : The dataset is an in-house clinical dataset with 74 general skin condition classes, which has more fine-grained classes built on the 9 classes of MMT-09. It includes general skin conditions like inflammatory, infective, benign proliferations, melanocytic, eczema. We chose the same 38,476 images as our evaluation dataset.
@@
Reader study1: Early melanoma detection based on SDDI-Alfred (2 classes) : The dataset consists of 179 serial dermoscopic imaging sequences from 122 patients, totaling 730 dermoscopic images. The patients were recruited from a private specialist dermatology clinic, with follow-up periods ranging from January 2007 to December 2019. Both melanoma and benign lesions that underwent short- or long-term sequential digital dermoscopic imaging (SDDI) at least once prior to biopsy were included. The dataset is well-balanced, with 90 benign lesions and 89 malignant lesions, encompassing both invasive and in situ melanomas. Of the 89 melanomas, 34 (38.2%) were invasive, with a mean Breslow thickness of 0.5 mm, while 55 (61.8%) were in situ. All lesions were monitored via digital dermoscopy, excised due to clinical concerns, and confirmed by pathological examination. The number of images per sequence varied from 1 to 12, with an average of approximately 4 images per sequence.
@@
Melanoma metastasis prediction and prognosis based on CombinMel dataset (2 or 3 classes) : This dataset encompasses 680 dermoscopic images of invasive melanoma from 370 patients recruited across 10 hospital sites in multiple countries, including Australia and 5 European nations. For large melanomas, multiple images were captured to ensure comprehensive coverage of the entire lesion area. Among the 370 melanoma cases, the majority 261 (70.5%) are classified as Stage I, with 45 (12.2%) in Stage II, 61 (16.5%) in Stage III, and a small fraction of 3 (0.8%) in Stage IV. This distribution provides a comprehensive representation of melanoma stages, with a focus on earlier stages. Regarding metastasis status, 248 (67.0%) of cases showed no metastasis, while 66 (17.8%) presented with metastasis at the time of diagnosis. Additionally, 56 (15.1%) of cases developed metastasis during the follow-up period.
@@
This dataset comprises 3D total body photography (TBP) images captured using the VECTRA WB360 system (Canfield Scientific Inc., Parsippany, NJ, USA). The system employs 92 cameras to simultaneously capture cross-polarised 2D images with standardized lighting within seconds, which are then merged to create a high-fidelity 3D avatar of each patient’s entire skin surface. From these 3D avatars, individual lesion tiles were exported for further analysis.
@@
MYM study was approved by the Metro South Health Human Research Ethics Committee on 21 April 2016 (approval number: HREC/16/QPAH/125). Ethics approval has also been obtained from the University of Queensland Human Research Ethics Committee (approval number: 2016000554), Queensland University of Technology Human Research Ethics Committee (approval number: 1600000515) and QIMR Berghofer (approval number: P2271). The HOP study has received Human Research Ethics Committee (HREC) approval from Metro South Health HREC (HREC/17/QPAH/816) and The University of Queensland HREC (2018000074). The ComBineMel dataset is part of the Computer biomarkers evaluation of invasive melanoma (ComBine Mel) study. The study was approved by the Alfred Hospital Ethics Committee on 08 August 2023 (approval number: HREC/98200/Alfred-2023). The study follows the National Statement on Ethical Conduct in Human Research (2007) protocols. SDDI2 dataset is approved by the Ethics Review Board of the Medical University of Vienna. MMT data study is part of a research agreement study with Monash eResearch Centre and was approved through the Monash University Human Research Ethics Committee (MUHREC). The NSSI dataset is part of the Brisbane Naevus Morphology Study, circa 2009-2014. The study followed the Declaration of Helsinki protocols and was approved by the Princess Alexandra Hospital human research ethics committee. The ACEMID_path study has received approval from the Alfred Hospital Ethics Committee (approval number: 746/23) to share data accrued for registered trial ACTRN12619001706167 (ACEMID) under the Metro South Human Research Committee protocol HREC/2019/QMS/57206 and the University of Queensland Human Research Ethics Committee protocol 2019003077. The SDDI_Alfred study has received approval from the Alfred Hospital Ethics Committee (approval number: 198/19) for use of sequential dermoscopic imaging data. Only de-identified retrospective data was used for research, without the active involvement of patients.

### v2

@@ Methods
MYM cohort (TBP). The MYM cohort [50] represents an in-house dataset designed to study the natural history of melanocytic naevi. This cohort, which included 193 participants from a population-based cohort in Australia, was recruited from the electoral roll, with eligibility criteria including. The cohort maintained a relatively balanced gender distribution with a slight female predominance (42.5% female, 57.5% male). The mean age of participants was 52.2 years (SD ± 12.0 years), with ages ranging from 23 to 70 years. The study population was predominantly of European/British ancestry (97.9%). Regarding sun response characteristics, 26.9% of participants reported burning easily and tanning slightly, 31.4% reported moderate burning and gradual tanning, and 11.0% reported rarely burning and tanning well, representing a predominantly fair-skinned population with varying degrees of UV sensitivity. 3D TBP was conducted using a VECTRA WB360 (Canfield Scientific Inc., Parsippany, NJ, USA), which instantaneously captures 92 cross-polarised 2D images with standardized lighting and subsequently merges them to create a 3D avatar. The average number of lesion tiles per subject from TBP was approximately 500. Demographic factors were collected using standard questionnaires, with clinical characteristics collected by research assistants. The final dataset of automatically detected lesion image tiles $\geq$ 2 mm in diameter comprises 405,856 images.
HOP cohort (TBP). TThe HOP study [49] is an in-house sequential dataset targeting high-risk melanoma individuals. The study cohort comprised 314 participants (120 males, 194 females) with a mean age of 56.1 years (SD = 12.8). The majority of participants were of European/British or other European ancestry (94.6%), with small proportions of other/mixed ancestries (5.4%). Most participants were born in Australia (85.4%), with 14.6% born overseas. In terms of skin cancer history, 96.8% of participants reported previous melanoma while 3.2% had no melanoma history. For non-melanoma skin cancer, 62.1% reported previous occurrence while 37.9% had no such history. Innate skin color (non-sun exposed skin) was reported as fair for 85.7%, medium for 14.0%, and olive for 0.3% of the cohort. Inclusion criteria for the HOP study were: at least one melanoma (including in situ) diagnosed before the age of 40 years, two or more melanomas (including in situ) diagnosed before the age of 65, a strong family history (two or more first-degree relatives affected) and/or known pathogenic mutations in a hereditary melanoma gene, and/or a diagnosis of dysplastic naevus syndrome. As with the MYM cohort, 3D TBP imaging was performed using the VECTRA WB360 system, following the same imaging protocol and data collection procedures. Demographic and clinical data were similarly collected through standardized questionnaires and by research assistants.
@@
ACEMID pathology pilot study. This dataset comprises 54 patients from two Australian centers: Queensland PAH (48.1%) and NSW MIA (51.9%), with a slight male predominance (57.4%) and ages ranging from 19 to 75 years (mean 53.4 years). Most patients (81.5%) were classified as ”Very High” risk for melanoma, with others being ”High” risk (14.8%) or ”Low/Average” risk (1.9%). Based on the first pathologist’s diagnoses, lesions were predominantly naevi (68.5%, including common/dermal/congenital and dysplastic variants), followed by melanomas (24.1%, mostly in situ), and other lesions (7.4%). While most patients (66.7%) had single lesions examined, others had multiple lesions reviewed (2-5 lesions per patient). Notable diagnostic variability between pathologists was observed, particularly in assessing dysplastic naevi and early melanomas.
NSSI. The NSSI dataset is an in-house sequential collection containing 29,832 dermoscopic images from 1,254 individuals residing in Brisbane, Australia. These images were collected as part of the Brisbane Naevus Morphology Study (2009-2014) using a digital dermatoscope attached to a Fotofinder ATBM imaging system, producing 768 × 576 pixel images at 96 dpi. The cohort comprised 48% male and 52% female participants, with a median age of 46 years (range: 11-88 years, $<$ 1% under 18). Just under half the cohort (47%) had a personal history of melanoma. Skin color was predominantly fair (51%, Fitzpatrick Type 1), followed by medium (41%, Type 2) and olive (8%, Type 3). The majority of participants reported British/Irish ancestry (87%), followed by Western/Northern European (21%), with smaller proportions of Southern European (3%), Asian (2%), and other ancestries ($<$ 1% each). The study design included up to 7 time points per participant at 6-month intervals over 3 years, with some variation due to enrollment timing and follow-up adherence. Individual lesions were assigned consistent identification numbers across visits to facilitate temporal analysis.
Edu1 & Edu2. The Edu1 and Edu2 datasets comprise 81,947 and 67,430 clinical images respectively, curated from in-house educational resources. These comprehensive collections cover a wide range of dermatological conditions including inflammatory and autoimmune disorders (such as psoriasis, and atopic dermatitis), common infections (such as herpes simplex, molluscum contagiosum, and tinea corporis), pigmentary disorders (like melasma and vitiligo), nail conditions (such as psoriatic nail disease and onychomycosis), vascular lesions (like port-wine stains and pyogenic granulomas), and both benign and malignant tumors (such as melanoma, basal cell carcinoma, and squamous cell carcinoma). The datasets also include rare conditions and genetic disorders.
@@
Sequential data preprocessing for lesion change detection. Our proposed sequential data preprocessing method consists of dark corner removal, skin inpainting, hair removal, image registration, and lesion segmentation. For the first two steps, we follow the approach outlined in [68]. Given an image with or without dark corner artifacts (DCA), we first convert it to grayscale. We then extract the contour by applying OpenCV’s [69] binary threshold function to the grayscale image, empirically setting the threshold at 100, and using the findContours function with RETR_TREE mode and CHAIN_APPROX_SIMPLE method. We identify the largest area contour in the image, which most closely matches the edge of the DCA, by calculating the area of all existing contours. Using OpenCV’s minEnclosingCircle function, we capture a circular area that encompasses this largest contour. To mitigate the effect of gradient colors at the dark corner edges, we scale this circle down to 80% of its original radius and convert it into a binary mask. Finally, we inpaint the original image using this mask, employing OpenCV’s implementation of the Telea algorithm with an inpaint radius of 10. Following dark corner removal and inpainting, we implement a hair removal step to further improve image quality and facilitate more accurate registration. This process begins by converting the image to grayscale. We then apply a black hat morphological operation using a 17$\times$17 structuring element to isolate dark, thin structures (hairs) from the background. The resulting image is thresholded to create a binary mask of the detected hair structures. Finally, we use OpenCV’s inpaint function with the Telea algorithm to fill in the hair regions, effectively removing them from the image. This hair removal step is crucial for improving the accuracy of subsequent image registration and analysis. For image registration, we implement an AKAZE [70] feature-based approach. The process begins by detecting key points and computing descriptors using the AKAZE algorithm, which is particularly effective for non-linear scale spaces. We utilize OpenCV’s AKAZE_create function, setting the descriptor size to 0, threshold to $9\times 10^{-5}$, and number of octaves to 4. Keypoint matching is performed using a Brute Force matcher with Hamming distance and cross-checking enabled. To refine the matches and estimate the transformation between images, we employ the RANSAC (Random Sample Consensus) algorithm implemented via skimage.measure.ransac. This estimates an EuclideanTransform model, which accounts for rotation and translation between the images. The resulting transformation is then applied to warp one image onto the other using skimage.transform.warp with reflection padding and linear interpolation.
@@
Early melanoma detection (Reader study 1). We fine-tuned our foundation model on the private SDDI-Alfred dataset[54] using a ten-fold cross-validation approach. We utilized cross-entropy loss with a learning rate of $5\times 10^{-4}$. We train models for 50 epochs with a warmup of 10 epochs. The model exhibiting the best AUROC on the validation set is selected as the final model. We then employed an out-of-fold (OOF) prediction approach to generate melanoma predictions for all sequential images. For each image sequence, we recorded the time point at which the model first made a correct diagnosis of melanoma; otherwise, the model was considered to have failed in detecting the melanoma. While biopsy serves as our reference standard, we aimed to explore the algorithm’s potential to detect early signs of melanoma progression. Our study focused on identifying suspicious changes in sequential images prior to clinical diagnosis, with the goal of enabling earlier intervention when melanomas are most treatable. For the human evaluation, 12 clinicians—seven dermatologists with over five years of experience and five dermatology residents with less than five years of experience—were invited to assess the serial dermoscopic data. The images were presented to the reviewers using Qualtrics™ (Provo, UT, USA), with the reviewers blinded to the true diagnoses. For each case, information such as the patient’s age, sex, lesion location, and date of imaging was provided. Initially, only the first dermoscopic image in the sequence was shown, and reviewers were asked to classify the lesion as either benign or malignant. As they progressed through the sequence, side-by-side image comparisons were made available to assess changes over time. Once a diagnosis was submitted, it could not be revised. To mitigate bias, we included 10 single time-point melanoma images, preventing reviewers from assuming that the first image in a series was benign. We then compared the diagnostic performance of the clinicians with our model, focusing on the time point at which a malignant diagnosis was first made by either the clinicians or the algorithm.
@@
Human-AI collaboration for skin cancer diagnosis (Reader study 2). The reader study was conducted using DermaChallenge, a web-based platform developed and hosted by the Medical University of Vienna for online education on dermatoscopy, as described in previous studies [75, 76]. To ensure proper authentication and data management, readers were required to register with a unique username, valid email address, and password. Active users on the platform, who previously actively agreed to be contacted, were recruited via a single email. Before commencing to the study phase, all users had to finish three introduction levels to be familiarized with the platforms’ user interface and image types. The number of correct answers in the first iteration of these levels normalized against the mean score of the entire DermaChallenge platform user base, served as a score of experience. Users were grouped into “Low” (n=11), “Medium” (n=21), and “High” (n=9) experience based on quantiles with cuts at 0.25 and 0.75 probability (R stats::quantile() function). Within the study level, users were shown batches of 10 images, randomly selected from a pool of 1,511 images, i.e. the ISIC 2018 Task 3 test set, with a predefined diagnosis distribution (AKIEC: 1, BCC: 1, BKL: 1, DF: 1, VASC: 1, MEL: 2, NV: 3). For each image a user had to choose one diagnosis out of seven options, and subsequently again after assistance from our foundation model, presented as multi-class probabilities visualized as bars and numbers for each class. Readers had the flexibility to complete multiple survey rounds with different image batches at their discretion, incompletely answered batches were omitted. The study was conducted online from August 20 to September 12, 2024, during which we collected data from 41 raters. Our foundation model for decision support utilized a weighted random sampler strategy, following the approach from [75] but excluding test-time augmentation. The model demonstrated robust performance, achieving an 80.4% mean (macro-averaged) recall, with notably high recall rates for critical skin lesions: 87.2% for melanoma and 86.0% for Basal Cell Carcinoma (BCC).
Human-AI collaboration for 128 skin conditions differential diagnosis (Reader study 3). The reader study was conducted using a web-based platform developed for online dermatological assessment. A total of 37 healthcare professionals participated in the study, categorized into two groups based on specialization: a dermatology group (n=20) comprising 9 dermatology specialists and 11 specialty trainees, and a generalist group (n=17) including 7 GPs, 7 general medicine practitioners, and 3 other healthcare professionals (nursing, clinical trial assistants) who manage skin conditions within their broader practice scope. This grouping strategy reflects the real-world clinical setting where non-dermatologist healthcare professionals routinely perform initial skin assessments. The diverse range of 128 skin conditions enabled evaluation of diagnostic performance between dermatologically trained professionals and those with general medical training. Readers were presented with clinical images and asked to provide their assessment through a structured questionnaire. Each participant rated image quality on a 5-point scale (from ”Not at all” to ”Completely” assessable), provided a primary diagnosis through free-text entry, and optionally listed two differential diagnoses ranked by likelihood. Diagnostic confidence was recorded on a 4-point scale (1: Not at all confident, 2: Somewhat confident, 3: Confident, 4: Highly confident). Following their initial assessment, readers were shown PanDerm’s top three predicted diagnoses and given the opportunity to maintain or modify their original diagnosis and differential diagnoses, followed by a reassessment of their confidence using the same 4-point scale. The study collected 1,342 responses between 07/01/2025 and 10/02/2025. Prior to evaluation, four experienced dermatologists collaboratively developed a standard ontology tree to systematically categorize the 128 skin conditions and facilitate expert evaluation (Extended Data Fig 5). The evaluation process involved multiple expert assessors who independently scored diagnostic accuracy using a 4-point scale: 4 (direct match with pre-defined term in the ontology), 3 (match within same diagnostic category in the ontology), 2 (inconsequential misdiagnosis), and 1 (significant mismatch, potentially dangerous misdiagnosis). To ensure robust assessment, each case was evaluated by three assessors, with cases showing significant scoring discordance (differences between 3/4 vs 1/2) reviewed in consensus meetings to establish final scores. For the top-3 accuracy evaluation, both human readers and AI assistance were evaluated based on whether the correct diagnosis appeared within their top three diagnostic choices.
@@
Evaluation metrics. For multi-class tasks, we primarily use a weighted F1 score, which averages class-specific F1 scores (harmonic means of precision and recall) weighted by class size. It addresses class imbalance in multi-class scenarios. For binary classification, we primarily use AUROC (Area Under the Receiver Operating Characteristic curve), measuring the model’s ability to distinguish between classes across all classification thresholds. An AUROC of 1.0 indicates perfect classification, while 0.5 suggests random guessing. This metric is particularly useful for imbalanced datasets and when we need to evaluate trade-offs between true positive and false positive rates. For the three reader studies, we report accuracy (top-1 or top-3). In skin lesion segmentation, we use the Dice Similarity Coefficient (DSC) and Jaccard index (JAC) to assess segmentation quality. For TBP-based melanoma screening, we primarily report the sensitivity (recall) in malignant lesions, focusing on the model’s ability to correctly identify malignant cases.
@@
DermNet [44] (23 classes) : DermNet contains 19,559 clinical images, the dataset consists of images of 23 types of skin diseases, captures common clinical presentations including inflammatory conditions (eczema, psoriasis), infections (bacterial, viral, fungal), and neoplastic diseases.
Fitzpatrick17K [61] (114 classes) : This dataset comprises 16,577 clinical images annotated with both dermatological diagnoses and Fitzpatrick skin types (I-VI). It encompasses 114 distinct conditions (minimum 53 images per condition) spanning major dermatological categories: inflammatory dermatoses (psoriasis, lichen planus, various eczematous conditions), cutaneous malignancies (melanoma, morpheiform and solid-cystic variants of BCC, SCC), papulosquamous disorders (pityriasis rosea, pityriasis rubra pilaris), autoimmune conditions (lupus erythematosus, bullous diseases), benign neoplasms (seborrheic keratosis, dermatofibroma), and various other clinically significant entities (acanthosis nigricans, granuloma annulare, necrobiosis lipoidica).
@@
MMT-74 (74 classes) : The MMT-74 dataset is a comprehensive in-house clinical collection comprising 38,476 dermatological images across 74 detailed skin condition classes, building upon and refining the broader 9-class structure of MMT-09. This structured dataset encompasses diverse dermatological conditions, including detailed classifications of basal cell carcinoma variants (nodular, pigmented, superficial, and recurrent), melanocytic lesions with specific pattern recognition (such as acral patterns and various nevus types), inflammatory disorders (dermatitis, psoriasis), benign proliferations (including seborrheic keratosis variants), and vascular lesions (angiomas, telangiectasias). The dataset was specifically designed to evaluate deep learning models’ performance across a diverse and clinically relevant range of skin conditions, with categories spanning inflammatory, infective, benign proliferations, melanocytic, and eczema classifications.
SD-128 (128 classes) : This dataset encompasses 5,619 clinical images covering 128 dermatological conditions spanning the complete spectrum of clinical practice. The dataset provides substantial coverage of inflammatory dermatoses, ranging from common presentations (such as psoriasis and atopic dermatitis) to less common entities (like leukocytoclastic vasculitis). It includes diverse infectious diseases of bacterial, viral, and fungal etiologies, as well as a comprehensive range of proliferative lesions from benign nevi to malignant melanomas. The collection also extends to appendageal disorders, physical trauma-related changes, nail disorders, and hair loss conditions. This extensive compilation represents both frequently encountered conditions in everyday practice and challenging rare cases, providing a robust resource for clinical diagnostic support. This dataset contains 5,619 clinical images encompassing diverse dermatological conditions commonly encountered in clinical practice. The dataset provides substantial coverage of inflammatory conditions from common presentations (psoriasis, atopic dermatitis) to less common entities (leukocytoclastic vasculitis), various infectious diseases spanning bacterial, viral, and fungal etiologies, a range of proliferative lesions from benign nevi to malignant melanomas, as well as appendageal disorders and physical trauma-related changes. We utilized 10% of the data stratified by disease labels for benchmark evaluation. Additionally, we selected 200 images stratified by disease classes for our reader study., which were further annotated with the Fitzpatrick skin tone scale (I-VI) by three expert dermatologists to enable subgroup performance analysis.
@@
Early melanoma detection based on SDDI-Alfred (2 classes) : The dataset consists of 179 serial dermoscopic imaging sequences from 122 patients, totaling 730 dermoscopic images. The patients were recruited from a private specialist dermatology clinic, with follow-up periods ranging from January 2007 to December 2019. The study population showed distinct characteristics between melanoma and benign groups: melanoma patients had a mean age of 56.6 years (SD=11.8) compared to 49.6 years (SD=11.4) in the benign group, with slightly different gender distributions (53.9% male in melanoma vs 40.0% male in benign cases). Both melanoma and benign lesions that underwent short- or long-term sequential digital dermoscopic imaging (SDDI) at least once prior to biopsy were included. The dataset is well-balanced, with 90 benign lesions and 89 malignant lesions. Of the 89 melanomas, 34 (38.2%) were invasive, with a mean Breslow thickness of 0.5 mm, while 55 (61.8%) were in situ. The melanoma subtypes included invasive SSM (36.0%), in situ SSM (31.4%), unspecified in situ (18.0%), lentigo maligna (12.3%), and invasive LMM (2.2%). The benign lesions were predominantly dysplastic naevi (40.0%), followed by compound naevi (27.8%), junctional naevi (18.9%), and intradermal naevi (8.9%). Anatomically, lesions were most commonly located on the lower limb (29.2% melanoma, 26.7% benign) and back (23.5% melanoma, 25.6% benign). All lesions were monitored via digital dermoscopy, excised due to clinical concerns, and confirmed by pathological examination. The number of images per sequence varied from 1 to 12, with an average of approximately 4 images per sequence.
@@
Melanoma metastasis prediction and prognosis based on CombinMel dataset (2 or 3 classes) : The CombinMel dataset encompasses 680 dermoscopic images of invasive melanoma from 370 patients recruited across 10 hospital sites in multiple countries, including Australia and 5 European nations. For large melanomas, multiple images were captured to ensure comprehensive coverage of the entire lesion area. The study population showed a relatively balanced distribution in both age (51.4% $>$60 years, 48.6% $\leq$60 years) and gender (47.3% female, 53.5% male). Regarding disease staging, the majority of cases were classified as Stage I (70.5%), followed by Stage III (16.5%), followed by Stage II (12.2%), and Stage IV (0.8%). In terms of T classification, T1a was the most common (59.2%), followed by T2a (18.6%) and T4b (13.2%). Sentinel lymph node biopsy (SLNB) was not performed in most cases (71.6%), with 10.8% positive and 17.6% negative results among those tested. For nodal status, N1 disease was the most common (10.8%), followed by N2 (3.8%) and N3 (1.8%). Regarding metastasis status, 248 (67.0%), of cases showed no metastasis, while 66 (17.8%) presented with metastasis at the time of diagnosis. Additionally, 56 (15.1%) of cases developed metastasis during the follow-up period.
@@
This dataset comprises 3D total body photography (TBP) images captured using the VECTRA WB360 system (Canfield Scientific Inc., Parsippany, NJ, USA). The system employs 92 cameras to simultaneously capture cross-polarised 2D images with standardized lighting within seconds, which are then merged to create a high-fidelity 3D avatar of each patient’s entire skin surface. From these 3D avatars, individual lesion tiles were exported for further analysis. Unlike standalone clinical photographs, TBP represents a higher-order imaging modality where 2D tiles are systematically derived from comprehensive 3D reconstructions, maintaining intrinsic spatial relationships within the 3D framework. The standardized acquisition process with calibrated lighting and positioning enables systematic capture of the entire body surface with overlapping views. This provides consistent anatomical landmarks and contextual information critical for comprehensive patient-level assessment, including evaluation of skin phenotype patterns, lesion measurements, and application of the ”ugly duckling” sign. The captured images undergo specific calibration and stitching processes, resulting in standardized 2D tiles that maintain consistent quality and perspective across all body regions.
@@
MYM study was approved by the Metro South Health Human Research Ethics Committee on 21 April 2016 (approval number: HREC/16/QPAH/125). Ethics approval has also been obtained from the University of Queensland Human Research Ethics Committee (approval number: 2016000554), Queensland University of Technology Human Research Ethics Committee (approval number: 1600000515) and QIMR Berghofer (approval number: P2271). The HOP study has received Human Research Ethics Committee (HREC) approval from Metro South Health HREC (HREC/17/QPAH/816) and The University of Queensland HREC (2018000074). The ComBineMel dataset is part of the Computer biomarkers evaluation of invasive melanoma (ComBine Mel) study. The study was approved by the Alfred Hospital Ethics Committee on 08 August 2023 (approval number: HREC/98200/Alfred-2023). The study follows the National Statement on Ethical Conduct in Human Research (2007) protocols. SDDI2 dataset is approved by the Ethics Review Board of the Medical University of Vienna. MMT data study is part of a research agreement study with Monash eResearch Centre and was approved through the Monash University Human Research Ethics Committee (MUHREC). The NSSI dataset is part of the Brisbane Naevus Morphology Study, circa 2009-2014. The study followed the Declaration of Helsinki protocols and was approved by the Princess Alexandra Hospital human research ethics committee. The ACEMID_path study has received approval from the Alfred Hospital Ethics Committee (approval number: 746/23) to share data accrued for registered trial ACTRN12619001706167 (ACEMID) under the Metro South Human Research Committee protocol HREC/2019/QMS/57206 and the University of Queensland Human Research Ethics Committee protocol 2019003077. The SDDI_Alfred study has received approval from the Alfred Hospital Ethics Committee (approval number: 198/19) for the use of sequential dermoscopic imaging data. Only de-identified retrospective data was used for research, without the active involvement of patients.